# Re-Envisioning Command and Control

### v1

@@ III-C Developing Adaptive Collective Knowledge Systems
As introduced in the SSG Hill data wrangler role and expanded in the LTC Mostert AAR example, C2 system personnel will be responsible for guiding and teaching the iCOAs-SA to effectively adapt throughout C2. We believe that this guidance will be necessary for two primary reasons: first, the nature of the C2 is so complex that humans will be needed to provide inputs for the systems to learn from (e.g., demonstrations, decisions, feedback) across the multiple steps of the C2 operations process as well as within different stages of data processing, integration, and interpretation. As discussed before, as the iCOAs-SA adapts, the exact nature and roles of these C2 personnel may shift; however, we propose that changing environments and mission requirements drive the need for continued human involvement across steps and stages. Second, humans will be critical to control mechanisms, such as human-machine AARs emplaced by trained leaders (see LTC Mostert), that modulate at what times and from which inputs the suite adapts to ensure effective, ethical, and moral behaviors over time. We expect that the changing nature of the human-machine interactions will be significant enough to influence how future Soldiers are recruited, assigned positions, and trained.
@@ IV Conclusions
While our approach largely focuses on providing examples of a C2 system possible in the far future, we posit that by focusing on human-machine partnerships as a core principle of system development, intelligent systems could be designed that intentionally evolve over time. That is, initially, suites of technologies could operate as aids in a largely human-dominated system akin to current C2, while over time with experience and refinement, the nature of the interactions could evolve to the proposed partnerships described herein. Further, we acknowledge that while we frame the problem in terms of human-machine partnerships, progress in human-centric and machine-centric approaches will be necessary and effective when integrated with the types of advancements envisioned herein.
@@ Acknowledgements
The authors would like to thank Benjamin T. Files, Nicholas Waytowich, Alfred Yu, Vernon Lawhern, and Steven Thurman for engaging in helpful discussions.

### v2

@@ III-C Developing Adaptive Collective Knowledge Systems
As introduced in the SSG Hill data wrangler role and expanded in the LTC Mostert AAR example, C2 system personnel will be responsible for guiding and teaching the iCOAs-SA to effectively adapt throughout C2. We believe that this guidance will be necessary for two primary reasons: first, the nature of the C2 is so complex and human-centric (e.g., see ethical, moral, legal, political considerations) that humans will need to provide inputs for the systems to learn from (e.g., demonstrations, decisions, feedback) across the multiple steps of the C2 operations process as well as within different stages of data processing, integration, and interpretation. As discussed before, as the iCOAs-SA adapts, the exact nature and roles of these C2 personnel may shift; however, we propose that changing environments and mission requirements drive the need for continued human involvement across steps and stages. Further, we assume that when iCOS-SA is introduced in novel environments or missions, the need for human guidance will be ramped up to identify and mitigate potentially ineffective behaviors before subsiding as the system learns. Second, humans will be critical to control mechanisms, such as human-machine AARs emplaced by trained leaders (see LTC Mostert), that modulate at what times and from which inputs the suite adapts to ensure effective, ethical, and moral behaviors over time. We expect that the changing nature of the human-machine interactions will be significant enough to influence how future Soldiers are recruited, assigned positions, and trained. To be effective, human-guided iCOAs-SA will have to be sensitive to differences between various interacting humans (e.g., different preferences or biases), as well as sensitive to human dynamics (e.g., state changes that may influence biases or errors generally).
@@ IV Conclusions
While our approach largely focuses on providing examples of a C2 system possible in the far future, we posit that by focusing on human-machine partnerships as a core principle of system development, intelligent systems could be designed that intentionally evolve over time. That is, initially, suites of technologies could operate as aids in a largely human-dominated system akin to current C2, while over time with experience and refinement, the nature of the interactions could evolve to the proposed partnerships described herein. Consistent with this, there are several recently-developed AI tools for the MDMP focused on accelerating mission analysis, such as understanding commander’s intent [17], and COA development and analysis [18, 15, 19]. Further, we acknowledge that while we frame the problem in terms of human-machine partnerships, progress in human- and machine-centric approaches will be necessary and effective when integrated with the types of advancements envisioned herein, particularly with respect to the technology and infrastructure available at the tactical edge. For instance, mirroring industry trends of integrating AI into devices such as smartphones and personal computers, we predict that the computing and power requirements for future AI technologies will be met for warfighters at the tactical edge.
@@ Acknowledgements
The authors would like to thank Benjamin T. Files, Nicholas Waytowich, Alfred Yu, Vernon Lawhern, Steven Thurman, Howie Brewington, and Jay Goodwin for engaging in helpful discussions.